OpenStudy (anonymous):

find the quotient. 4x^3y+12x^2y^3-20xy divided by or over 4xy

OpenStudy (anonymous):

@sleepyjess

OpenStudy (igreen):

Divide all the whole numbers by 4.

OpenStudy (igreen):

Subtract all the exponents with a variable of 'x' by 1, and subtract all the exponents with a variable of 'y' by 1.

OpenStudy (igreen):

\(\dfrac{4}{4}x^{3-1}y^{1-1} + \dfrac{12}{4}x^{2-1}y^{3-1} - \dfrac{20}{4}x^{1 - 1}y^{1-1}\) Simplify @LOUISndGUCCI

OpenStudy (igreen):

Got it? @LOUISndGUCCI

OpenStudy (anonymous):

yes i got it thank you

OpenStudy (igreen):

No problem.